# Solution of some tiling open problems of Propp, Lai, and some related results

### v2

@@ 4.3. Two families of polynomials and their connection to the dented regions
((((a)))). For $k$ even, we have
$[\widehat{H^{\prime}}_{\mathbf{l}}(a,b,k)^{+},\widehat{H^{\prime}}_{\mathbf{l}}(a,b,k)^{-}]=\begin{cases}[R_{\emptyset,{\mathbf{l}}}((a+k-3)/2),\bar{R}_{{\mathbf{l}}-1,\emptyset}((a+1)/2)],&\text{$a$ odd, $l_{1}=1$,}\\ [R_{\emptyset,{\mathbf{l}}}((a+k-3)/2),R_{{\mathbf{l}}-1,\emptyset}((a+1)/2)],&\text{$a$ odd, $l_{1}>1$,}\\ [R_{\emptyset,{\mathbf{l}}^{(m)}}((a+k-2)/2),\bar{R}_{{\mathbf{l}},\emptyset}(a/2)],&\text{$a$ even.}\end{cases}$ (4.14)
((((b)))). For $k$ odd, we have
$[\widehat{H^{\prime}}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{+},\widehat{H^{\prime}}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{-}]=\begin{dcases}[\bar{R}_{{\mathbf{l}},{\mathbf{q}}}((a+k-2)/2),{R}_{{\mathbf{q}},{\mathbf{l}}^{(m)}}((a+1)/2)],&\text{$a$ odd,}\\ [\bar{R}_{{\mathbf{l}},{\mathbf{q}}^{(n)}}((a+k-1)/2),{R}_{{\mathbf{q}},{\mathbf{l}}}(a/2)],&\text{$a$ even,}\end{dcases}$ (4.15)
@@
$[\widehat{\bar{H}^{\prime}}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{+},\widehat{\bar{H}^{\prime}}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{-}]=\begin{dcases}[{R}_{{\mathbf{l}},{\mathbf{q}}}((a+k-2)/2),\bar{R}_{{\mathbf{q}},{\mathbf{l}}^{(m)}}((a+1)/2)],&\text{$a$ odd,}\\ [{R}_{{\mathbf{l}},{\mathbf{q}}^{(n)}}((a+k-1)/2),\bar{R}_{{\mathbf{q}},{\mathbf{l}}}(a/2)],&\text{$a$ even.}\end{dcases}$ (4.16)
@@
Consider one of the regions $H^{\prime}_{\mathbf{l}}(a,b,k)$, $H^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ or $\bar{H}^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$, and denote it for short by $H^{\prime}(a)$. Denote by $H(a)$ the corresponding unprimed region of the same parameters ((((the number of lozenge tilings of $H(a)$ is given by Theorem 4.2 and the formulas in Theorem 4.1)))). Then we have
@@
Remark $4$. The above result has the following pleasing geometrical interpretation. Suppose we place a two-sided mirror along the axis $\ell^{\prime}$ with respect to which the holes in $H^{\prime}(a)$ are symmetric. Consider the region seen in the mirror from its left (this is precisely $H(a+1)$), and also the region seen in the mirror from its right (which is precisely $H(a-1)$). Then, up to a simple factor which depends on the parity of $a$, $\operatorname{M}(H^{\prime}(a))$ is just the geometric mean of the numbers of tilings of these two regions.
@@
For lists of strictly increasing positive integers ${\mathbf{l}}=(l_{1},\ldots,l_{m})$ and ${\mathbf{q}}=(q_{1},\ldots,q_{n})$, we have212121 When $m=0$ or $n=0$, the same conventions apply as for equations (4.1) and (4.2).
@@
[…]
@@
[…]
@@
This application of Theorem 2.2 involves two different zigzag cuts (as described before the statement of Theorem 2.2) and produces four half-regions $H^{+}$, $H^{+}$, $\widehat{H}^{+}$ and $\widehat{H}^{-}$. It is not hard to see that all these four belong to the two families defined in subsection 4.2. More precisely, we have
@@ 5. Nearly symmetric hexagons with a shamrock or fern hole
In this section we solve another open problem of Lai (namely, [24, Problem 27]), which asks to find and prove a simple product formula for the number of lozenge tilings of the region obtained from a symmetric $S$-cored hexagon by translating the shamrock hole one unit southeast; we call such a region a nearly symmetric $S$-cored hexagon (see Theorem 5.1). In addition, we prove a similar formula for $F$-cored hexagons — regions obtained from hexagons by removing a fern232323 Ferns were introduced in [6], where a simple product formula was proved for the number of tilings of a hexagon with a fern removed from its center. The number of tilings of an $F$-cored hexagon which is symmetric about a vertical axis was determined in [23]. — the union of contiguous triangular lobes of alternating orientation lined up along a horizontal lattice line — instead of a shamrock (such a region will be called a nearly symmetric $F$-cored hexagon; see Theorem 5.3).
@@
Let $H^{\prime}$ be a nearly symmetric $S$-cored hexagon in which the side-lengths of the triangular lobes are $\alpha$ for the top one, $\beta$ for the two bottom ones, and $\mu$ for the central one; let the side-lengths of the outer hexagon be $x+\alpha+2\beta$, $y+\mu$, $y+\alpha+2\beta$, $x+\mu$, $y+\alpha+2\beta$, $y+\mu$ ((((clockwise from top; see the picture on the left in Figure 25 for an example)))). Let $\ell^{\prime}$ be the vertical symmetry axis of the removed shamrock.
Denote by $H_{1}$ the region consisting of the union of the portion of $H^{\prime}$ to the right of $\ell^{\prime}$ with its reflection across $\ell^{\prime}$ ((((i.e., the region seen if $\ell^{\prime}$ was a mirror and one looks in it from the right; see the picture on the right in Figure 25 for an illustration)))); let $H_{2}$ be the analogous region determined by the portion of $H^{\prime}$ to the left of $\ell^{\prime}$. Then we have
@@
$f=\begin{dcases}\frac{(x+y+\alpha+2\beta+\mu)^{2}}{(x+\alpha+2\beta+\mu)(x+2y+\alpha+2\beta+\mu)},&\text{$x$ odd},\\ \frac{(x+y+\alpha+2\beta+\mu)^{2}}{x(x+2y+2\alpha+4\beta+2\mu)},&\text{$x$ even}.\\ \end{dcases}$ (5.2)
@@
We recall the “bowtie squeezing theorem” [10, Theorem 1]. A triad hexagon is a region obtained from a hexagon by removing three bowtie-shaped holes positioned as illustrated in [10, Figure 1]. Bowtie squeezing is an operation that turns a triad hexagon into another triad hexagon by squeezing part (or all) of a lobe of a bowtie into the other lobe; the position of the node of the squeezed bowtie as well as the shape of the other two bowties remain the same, but the relative position of the bowties and the boundary hexagon change (an illustrative example is shown [10, Figure 1] — see that paper for the details; the picture on the left in Figure 25 shows a special case). Then [10, Theorem 1] states that if the region $Q$ is obtained from the triad hexagon $R$ by a sequence of bowtie squeezing operations, then we have242424 Here $A$, $B$ and $C$ (resp., $A_{1}$, $B_{1}$ and $C_{1}$) are the nodes — i.e., the points where the two lobes meet — of the bowties in the triad hexagon $R$ (resp., $Q$), counterclockwise from top.
@@
[…]
@@
$\frac{\operatorname{M}(R)}{\operatorname{M}(Q)}=\dfrac{\operatorname{w}^{(R)}\dfrac{\prod_{\text{$A$ node of $R$}}\,\operatorname{k}_{A}^{(R)}}{\prod_{\text{$L$ support line of $R$}}\,\operatorname{k}_{L}^{(R)}}}{\operatorname{w}^{(Q)}\dfrac{\prod_{\text{$A$ node of $Q$}}\,\operatorname{k}_{A}^{(Q)}}{\prod_{\text{$L$ support line of $Q$}}\,\operatorname{k}_{L}^{(Q)}}}.$ (5.13)
@@
where $X$ (resp., $Y$) is the set of labels252525 The unit segments on $l$ inside $\overline{R}$ are labeled $1,2,3,\dotsc$ from left to right. of the up-pointing (resp., down-pointing) unit triangles with bases on $L$ which are removed from the outer hexagon of $\overline{R}$, while $X^{\prime}$ and $Y^{\prime}$ are the corresponding sets of labels for $\overline{Q}$.
@@
Let $H^{\prime}$ be a nearly symmetric $F$-cored hexagon in which the side-lengths of the triangular lobes of the removed ferns are $\alpha_{1},\alpha_{2},\cdots$ for the down-pointing ones, and $\beta_{1},\beta_{2},\cdots$ for the up-pointing ones. Set $\alpha=\alpha_{1}+\alpha_{2}+\cdots$ and $\beta=\beta_{1}+\beta_{2}+\cdots$, and let the side-lengths of the outer hexagon be $x+\alpha$, $y+\beta$, $y+\alpha$, $x+\beta$, $y+\alpha$, $y+\beta$ ((((clockwise from top)))). Let $\ell^{\prime}$ be the vertical symmetry axis of the removed fern.
@@
$f=\begin{dcases}\frac{(x+y+\alpha+\beta)^{2}}{(x+\alpha+\beta)(x+2y+\alpha+\beta)},&\text{$x$ odd},\\ \frac{(x+y+\alpha+\beta)^{2}}{x(x+2y+2\alpha+2\beta)},&\text{$x$ even}.\\ \end{dcases}$ (5.16)
@@ 6. Nearly symmetric hexagons with a lozenge hole
Using Theorem 2.2, we can also obtain a counterpart of a result of Fulmek and Krattenthaler which was presented in [17]. In [17] and [18], these authors considered the problem of enumerating lozenge tilings of a hexagon that has both vertical and horizontal symmetry — in this section, we will call this a symmetric hexagon262626 Doubly-symmetric would be more descriptive, but then naming the region obtained by translating the lozenge-shaped hole off the symmetry axis would become cumbersome. — which contain a fixed lozenge (or equivalently, tilings of a hexagon with a lozenge-shaped hole). In [17], they considered a symmetric hexagon, fixed a unit lozenge on the vertical [^27] Unlike in the current paper, in [17], the authors use a triangular lattice such that one family of lattice line is vertical. Thus, a horizontal symmetry axis in [17] corresponds to a vertical symmetry axis in the current paper. symmetry axis in an arbitrary position, and proved the following two results (Figure 27 illustrates the two types of regions considered).
@@
Let $m$ be a non-negative integer and $N$ a positive integer. The number of lozenge tilings of a hexagon with sides $N,2m,N,N,2m,N$, which contain the $l$-th lozenge on the symmetry axis which cuts through the sides of length $2m$, equals282828 The reason the formula is written as a prefactor times the shown triple product is that, by a classical result of MacMahon [27], the number of lozenge tilings of a hexagon of side-lengths $a$, $b$, $c$, $a$, $b$, $c$ equals $\prod_{i=1}^{a}\prod_{j=1}^{b}\prod_{k=1}^{c}\frac{i+j+k-1}{i+j+k-2}$.
@@
Let $m$ and $N$ be positive integers. The number of lozenge tilings of a hexagon with sides $N+1,2m-1,N+1,N+1,2m-1,N+1$, which contain the $l$-th lozenge on the symmetry axis which cuts through the sides of length $12m-$ 1, equals
@@
Let $n$ and $m$ be positive integers. The number of lozenge tilings of a hexagon with side lengths $2n,2m,2n,2n,2m,2n$ ((((clockwise from top left)))), which contain the lozenge just to the right of the center of the hexagon, equals
@@
Let $n$ be a non-negative integer and $m$ a positive integer. The number of lozenge tilings of a hexagon with side lengths $12n+1,2m-1,2n+1,2n+1,2m-1,2n+$ 1 ((((clockwise from top left)))), which contain the lozenge just to the right of the center of the hexagon, equals
@@
Similarly, for positive integers $m$ and $N$, let $V(N+1,2m-1,l)$ be the region obtained from the symmetric hexagon of side-lengths $12m-1,N+1,N+1,2m-1,N+1,N+$ 1 (clockwise from top) by deleting the $l$-th lozenge on the vertical symmetry axis. Similarly, let $V^{\prime}(N+1,2m-1,l)$ be the region obtained from the same hexagon by deleting the $l$-th lozenge on the vertical axis that is half a unit to the right of the vertical symmetry axis.
The first two results of Fulmek and Krattenthaler stated above (Theorem 6.1 and 6.2) are about the number of lozenge tilings of the regions $V(N,2m,l)$ and $V(N+1,2m-1,l)$. The main results of this section (see Theorems 6.6 and 6.8 below) give formulas for the number of lozenge tilings of the regions $V^{\prime}(N,2m,l)$ and $V^{\prime}(N+1,2m-1,l)$, and thus extends the results of Fulmek and Krattenthaler quoted as Theorems 6.3 and 6.4 by lending them the generality of the context of their results quoted as Theorems 6.1 and 6.2.
@@
Let $m$ be a positive integer and $N$ be a positive integer greater than $1$. The number of lozenge tilings of the region $V^{\prime}(N,2m,l)$ obtained from the hexagon with sides $2m,N,N,2m,N,N$ ((((clockwise from top)))) by deleting the $l$-th lozenge from the top on the vertical axis half a unit away from the vertical symmetry axis, is given by
@@
Let $m$ be a non-negative integer and $N$ be a positive integer greater than $1$. The number of lozenge tilings of the region $V^{\prime}(N-1,2m+1,l)$ obtained from the hexagon of sides $12m+1,N-1,N-1,2m+1,N-1,N-$ 1 ((((clockwise from top)))) by deleting the $l$-th lozenge from the top on the vertical axis half a unit away from the vertical symmetry axis, is given by
@@
For non-negative integers $m$ and $N$, let $S(N,m)$ be the pentagonal region depicted on the left in Figure 30; note that this region was previously denoted by $\overline{R}_{(1,\ldots,N),\emptyset}(m)$ in subsection 4.2. Also, for a non-negative integer $m$ and positive integers $l$ and $N$ such that $l\leq N$, we define the region $C(N,m,l)$ as follows: on $S(N,m)$, we label the $N$ rightmost vertical lozenges by $1,\ldots,N$, from top to bottom. Then we remove the vertical lozenge labeled by $l$ and assign the weight $\frac{1}{2}$ to the remaining $N-1$ vertical lozenges (see the picture on the right in Figure 30). In [17], Fulmek and Krattenthaler determined the number of lozenge tilings for $S(N,m)$ and the weighted count of lozenge tilings [^29] Our formula looks somewhat different compared to the one provided in [17] since we made some simplifications in the latter. of $C(N,m,l)$.
@@
((2.4)((2.4)( ( 2.4 ) and $(2.6)$ in [17])))) For any non-negative integer $m$ and positive integer $N$,
@@
((2.5)((2.5)( ( 2.5 ) and $(2.7)$ in [17])))) For any non-negative integer $m$ and positive integers $l$ and $N$ such that $l\leq N$,
@@
[Proof of Theorem 6.6] We apply Theorem 2.2 to the region $V^{\prime}(N,2m,l)$. To do that, we split the region using the two different zigzag lines along the vertical axis through the hole. While one of the zigzag lines splits the region into $S(N,m)$ and $C(N-1,m,l)$, the other splits it into $S(N,m-1)$ and $C(N-1,m+1,l)$ (see the two pictures in Figure 31). Thus, by Theorem 2.2, we have
@@
[Proof of Theorem 6.8] Theorem 2.2 is not applicable when $l=1$, because the region $R_{0}$ is not well-defined. Because of this, we consider the case $l=1$ separately. In this case, the region $V^{\prime}(N-1,2m+1,1)$ can be regarded as a trapezoid region with dents on its base (see Figure 32 and the explanation in its caption). The number of lozenge tilings of such regions was given in [11, Proposition 2.1] and one can check that the $l=1$ specialization of our claimed formula agrees with the expression obtained using [11, Proposition 2.1]. Assume therefore that $l\geq 2$, and apply Theorem 2.2. To do that, we again consider the two zigzag lines along the vertical symmetry axis of the lozenge hole. One of them divides the region into $S(N-2,m)$ and $C(N-1,m+1,l)$, while the other divides it into $S(N-2,m+1)$ and $C(N-1,m,l)$ (see Figure 33. Thus, by Theorem 2.2, we obtain

### v3

@@ 4.3. Two families of polynomials and their connection to the dented regions
((a)). For $k$ even, we have
[…]
((b)). For $k$ odd, we have
[…]
@@
[…]
@@
Consider one of the regions $H^{\prime}_{\mathbf{l}}(a,b,k)$, $H^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ or $\bar{H}^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$, and denote it for short by $H^{\prime}(a)$. Denote by $H(a)$ the corresponding unprimed region of the same parameters ((the number of lozenge tilings of $H(a)$ is given by Theorem 4.2 and the formulas in Theorem 4.1)). Then we have
@@
Remark $4$. The above result has the following pleasing geometrical interpretation. Suppose we place a two-sided mirror along the axis $\ell^{\prime}$ with respect to which the holes in $H^{\prime}(a)$ are symmetric. Consider the region seen in the mirror from its left (this is precisely $H(a+1)$), and also the region seen in the mirror from its right (which is precisely H(a1)H(a-1)). Then, up to a simple factor which depends on the parity of $a$, $\operatorname{M}(H^{\prime}(a))$ is just the geometric mean of the numbers of tilings of these two regions.
@@
For lists of strictly increasing positive integers ${\mathbf{l}}=(l_{1},\ldots,l_{m})$ and ${\mathbf{q}}=(q_{1},\ldots,q_{n})$, we have232323 When $m=0$ or $n=0$, the same conventions apply as for equations (4.1) and (4.2).
@@
$=\leavevmode\nobreak$ $\frac{B_{m,n}(x+l_{m}-m+1)}{B_{m,n}(x+l_{m}-m)}\cdot\prod_{i=1}^{m}\Bigg{[}\frac{(x+l_{m}-i+1)(x+l_{m}-m+n+l_{i}+2)}{(x+l_{m}-l_{i}+1)(x+l_{m}-m+n+i+2)}\Bigg{]}$
@@
$=\leavevmode\nobreak$ $\frac{(x+n+3)_{m}}{(x+n+1)_{m}}\cdot\frac{\langle x+3,x+n+1\rangle\langle x+\frac{5}{2},x+\frac{2n+3}{2}\rangle}{\langle x+2,x+n\rangle\langle x+\frac{3}{2},x+\frac{2n+1}{2}\rangle}$
@@
This application of Theorem 2.2 involves two different zigzag cuts (as described before the statement of Theorem 2.2) and produces four half-regions $^{+}, H+H^{+}$, $\widehat{H}^{+}$ and $\widehat{H}^{-}$. It is not hard to see that all these four belong to the two families defined in subsection 4.2. More precisely, we have
@@ 5. Nearly symmetric hexagons with a shamrock or fern hole
In this section we solve another open problem of Lai (namely, [24, Problem 27]), which asks to find and prove a simple product formula for the number of lozenge tilings of the region obtained from a symmetric $S$-cored hexagon by translating the shamrock hole one unit southeast; we call such a region a nearly symmetric $S$-cored hexagon (see Theorem 5.1). In addition, we prove a similar formula for $F$-cored hexagons — regions obtained from hexagons by removing a fern252525 Ferns were introduced in [6], where a simple product formula was proved for the number of tilings of a hexagon with a fern removed from its center. The number of tilings of an $F$-cored hexagon which is symmetric about a vertical axis was determined in [23]. — the union of contiguous triangular lobes of alternating orientation lined up along a horizontal lattice line — instead of a shamrock (such a region will be called a nearly symmetric $F$-cored hexagon; see Theorem 5.3).
@@
Let $H^{\prime}$ be a nearly symmetric $S$-cored hexagon in which the side-lengths of the triangular lobes are $\alpha$ for the top one, $\beta$ for the two bottom ones, and $\mu$ for the central one; let the side-lengths of the outer hexagon be $x+\alpha+2\beta$, $y+\mu$, $y+\alpha+2\beta$, $x+\mu$, $y+\alpha+2\beta$, $y+\mu$ ((clockwise from top; see the picture on the left in Figure 27 for an example)). Let $\ell^{\prime}$ be the vertical symmetry axis of the removed shamrock.
Denote by $H_{1}$ the region consisting of the union of the portion of $H^{\prime}$ to the right of $\ell^{\prime}$ with its reflection across $\ell^{\prime}$ ((i.e., the region seen if $\ell^{\prime}$ was a mirror and one looks in it from the right; see the picture on the right in Figure 27 for an illustration)); let $H_{2}$ be the analogous region determined by the portion of $H^{\prime}$ to the left of $\ell^{\prime}$. Then we have
@@
[…]
@@
We recall the “bowtie squeezing theorem” [10, Theorem 1]. A triad hexagon is a region obtained from a hexagon by removing three bowtie-shaped holes positioned as illustrated in [10, Figure 1]. Bowtie squeezing is an operation that turns a triad hexagon into another triad hexagon by squeezing part (or all) of a lobe of a bowtie into the other lobe; the position of the node of the squeezed bowtie as well as the shape of the other two bowties remain the same, but the relative position of the bowties and the boundary hexagon change (an illustrative example is shown [10, Figure 1] — see that paper for the details; the picture on the left in Figure 27 shows a special case). Then [10, Theorem 1] states that if the region $Q$ is obtained from the triad hexagon $R$ by a sequence of bowtie squeezing operations, then we have262626 Here $A$, $B$ and $C$ (resp., $A_{1}$, $B_{1}$ and $C_{1}$) are the nodes — i.e., the points where the two lobes meet — of the bowties in the triad hexagon $R$ (resp., $Q$), counterclockwise from top.
@@
where the weight $\operatorname{w}^{(R)}$ depends only on the shapes of the three bowties and their relative positions (the precise expression is given in [10, (2.2)], but we don’t need it here), and the point couples like $\operatorname{k}_{A}^{(R)}$ (resp., the line couples like $\operatorname{k}_{BC}^{(R)}$) have the following common definition: if $\operatorname{h}(n)$ is the hyperfactorial $\operatorname{h}(n)=0!\,1!\,\cdots\,(n-1)!$, the point couple $\operatorname{k}_{A}^{(R)}$ is defined to be $\operatorname{h}(u)\operatorname{h}(d)$, where $u$ and $d$ are the distances from $A$ to the pair of opposite sides of the outer hexagon towards which the $60^{\circ}$ degree double cone at $A$ contained in the bowtie points (for line couples, $u$ and $d$ are the distances from the line to the sides of the hexagon parallel to it).
@@
[…]
@@
where $X$ (resp., $Y$) is the set of labels272727 The unit segments on $l$ inside $\overline{R}$ are labeled $1,2,3,\dotsc$ from left to right. of the up-pointing (resp., down-pointing) unit triangles with bases on $L$ which are removed from the outer hexagon of $\overline{R}$, while $X^{\prime}$ and $Y^{\prime}$ are the corresponding sets of labels for $\overline{Q}$.
@@
Let $H^{\prime}$ be a nearly symmetric $F$-cored hexagon in which the side-lengths of the triangular lobes of the removed ferns are $\alpha_{1},\alpha_{2},\cdots$ for the down-pointing ones, and $\beta_{1},\beta_{2},\cdots$ for the up-pointing ones. Set $\alpha=\alpha_{1}+\alpha_{2}+\cdots$ and $\beta=\beta_{1}+\beta_{2}+\cdots$, and let the side-lengths of the outer hexagon be $x+\alpha$, $y+\beta$, $y+\alpha$, $x+\beta$, $y+\alpha$, $y+\beta$ ((clockwise from top)). Let $\ell^{\prime}$ be the vertical symmetry axis of the removed fern.
@@
[…]
@@ 6. Nearly symmetric hexagons with a lozenge hole
Using Theorem 2.2, we can also obtain a counterpart of a result of Fulmek and Krattenthaler which was presented in [17]. In [17] and [18], these authors considered the problem of enumerating lozenge tilings of a hexagon that has both vertical and horizontal symmetry — in this section, we will call this a symmetric hexagon282828 Doubly-symmetric would be more descriptive, but then naming the region obtained by translating the lozenge-shaped hole off the symmetry axis would become cumbersome. — which contain a fixed lozenge (or equivalently, tilings of a hexagon with a lozenge-shaped hole). In [17], they considered a symmetric hexagon, fixed a unit lozenge on the vertical [^29] Unlike in the current paper, in [17], the authors use a triangular lattice such that one family of lattice line is vertical. Thus, a horizontal symmetry axis in [17] corresponds to a vertical symmetry axis in the current paper. symmetry axis in an arbitrary position, and proved the following two results (Figure 29 illustrates the two types of regions considered).
@@
Let $m$ be a non-negative integer and $N$ a positive integer. The number of lozenge tilings of a hexagon with sides $N,2m,N,N,2m,N$, which contain the $l$-th lozenge on the symmetry axis which cuts through the sides of length $2m$, equals303030 The reason the formula is written as a prefactor times the shown triple product is that, by a classical result of MacMahon [27], the number of lozenge tilings of a hexagon of side-lengths $a$, $b$, $c$, $a$, $b$, $c$ equals $\prod_{i=1}^{a}\prod_{j=1}^{b}\prod_{k=1}^{c}\frac{i+j+k-1}{i+j+k-2}$.
@@
Let $m$ and $N$ be positive integers. The number of lozenge tilings of a hexagon with sides N+1,2m1,N+1,N+1,2m1,N+1N+1,2m-1,N+1,N+1,2m-1,N+1, which contain the $l$-th lozenge on the symmetry axis which cuts through the sides of length 2m12m-1, equals
@@
Let $n$ and $m$ be positive integers. The number of lozenge tilings of a hexagon with side lengths $2n,2m,2n,2n,2m,2n$ ((clockwise from top left)), which contain the lozenge just to the right of the center of the hexagon, equals
@@
Let $n$ be a non-negative integer and $m$ a positive integer. The number of lozenge tilings of a hexagon with side lengths 2n+1,2m1,2n+1,2n+1,2m1,2n+12n+1,2m-1,2n+1,2n+1,2m-1,2n+1 ((clockwise from top left)), which contain the lozenge just to the right of the center of the hexagon, equals
@@
Similarly, for positive integers $m$ and $N$, let V(N+1,2m1,l)V(N+1,2m-1,l) be the region obtained from the symmetric hexagon of side-lengths 2m1,N+1,N+1,2m1,N+1,N+12m-1,N+1,N+1,2m-1,N+1,N+1 (clockwise from top) by deleting the $l$-th lozenge on the vertical symmetry axis. Similarly, let $V^{\prime}(N+1,2m-1,l)$ be the region obtained from the same hexagon by deleting the $l$-th lozenge on the vertical axis that is half a unit to the right of the vertical symmetry axis.
The first two results of Fulmek and Krattenthaler stated above (Theorem 6.1 and 6.2) are about the number of lozenge tilings of the regions $V(N,2m,l)$ and V(N+1,2m1,l)V(N+1,2m-1,l). The main results of this section (see Theorems 6.6 and 6.8 below) give formulas for the number of lozenge tilings of the regions $V^{\prime}(N,2m,l)$ and $V^{\prime}(N+1,2m-1,l)$, and thus extends the results of Fulmek and Krattenthaler quoted as Theorems 6.3 and 6.4 by lending them the generality of the context of their results quoted as Theorems 6.1 and 6.2.
@@
Let $m$ be a positive integer and $N$ be a positive integer greater than $1$. The number of lozenge tilings of the region $V^{\prime}(N,2m,l)$ obtained from the hexagon with sides $2m,N,N,2m,N,N$ ((clockwise from top)) by deleting the $l$-th lozenge from the top on the vertical axis half a unit away from the vertical symmetry axis, is given by
@@
Let $m$ be a non-negative integer and $N$ be a positive integer greater than $1$. The number of lozenge tilings of the region $V^{\prime}(N-1,2m+1,l)$ obtained from the hexagon of sides 2m+1,N1,N1,2m+1,N1,N12m+1,N-1,N-1,2m+1,N-1,N-1 ((clockwise from top)) by deleting the $l$-th lozenge from the top on the vertical axis half a unit away from the vertical symmetry axis, is given by
@@
For non-negative integers $m$ and $N$, let $S(N,m)$ be the pentagonal region depicted on the left in Figure 32; note that this region was previously denoted by $\overline{R}_{(1,\ldots,N),\emptyset}(m)$ in subsection 4.2. Also, for a non-negative integer $m$ and positive integers $l$ and $N$ such that $l\leq N$, we define the region $C(N,m,l)$ as follows: on $S(N,m)$, we label the $N$ rightmost vertical lozenges by $1,\ldots,N$, from top to bottom. Then we remove the vertical lozenge labeled by $l$ and assign the weight $\frac{1}{2}$ to the remaining N1N-1 vertical lozenges (see the picture on the right in Figure 32). In [17], Fulmek and Krattenthaler determined the number of lozenge tilings for $S(N,m)$ and the weighted count of lozenge tilings [^31] Our formula looks somewhat different compared to the one provided in [17] since we made some simplifications in the latter. of $C(N,m,l)$.
@@
$((2.4)$ and $(2.6)$ in [17])) For any non-negative integer $m$ and positive integer $N$,
@@
$((2.5)$ and $(2.7)$ in [17])) For any non-negative integer $m$ and positive integers $l$ and $N$ such that $l\leq N$,
@@
[Proof of Theorem 6.6] We apply Theorem 2.2 to the region $V^{\prime}(N,2m,l)$. To do that, we split the region using the two different zigzag lines along the vertical axis through the hole. While one of the zigzag lines splits the region into $S(N,m)$ and C(N1,m,l)C(N-1,m,l), the other splits it into S(N,m1)S(N,m-1) and C(N1,m+1,l)C(N-1,m+1,l) (see the two pictures in Figure 33). Thus, by Theorem 2.2, we have
@@
[Proof of Theorem 6.8] Theorem 2.2 is not applicable when $l=1$, because the region $R_{0}$ is not well-defined. Because of this, we consider the case $l=1$ separately. In this case, the region $V^{\prime}(N-1,2m+1,1)$ can be regarded as a trapezoid region with dents on its base (see Figure 34 and the explanation in its caption). The number of lozenge tilings of such regions was given in [11, Proposition 2.1] and one can check that the $l=1$ specialization of our claimed formula agrees with the expression obtained using [11, Proposition 2.1]. Assume therefore that $l\geq 2$, and apply Theorem 2.2. To do that, we again consider the two zigzag lines along the vertical symmetry axis of the lozenge hole. One of them divides the region into S(N2,m)S(N-2,m) and C(N1,m+1,l)C(N-1,m+1,l), while the other divides it into S(N2,m+1)S(N-2,m+1) and C(N1,m,l)C(N-1,m,l) (see Figure 35. Thus, by Theorem 2.2, we obtain